Below are five comparable companies that compete in the Widget industry: Company V, Company W Company X, Company Y, and Company Z. Assume the greater the assets on the balance sheet, the larger the company. Use the financial statement information in the table below to help answer questions 1-6 Debt/Equity Total Aspets Tax Rate Net Income Company V Company W Company X 060 4.0000 100 1. Compared to larger companies in the widget industry, smaller companies in the Widget industry A. have a higher payout ratio B. have a lower retention ratio C. have a lower effective tax rote D. have a faster long term growth rate E. have more debt as a percentage of the capital structure 2. The CAPM was used to calculate the cost of equity (Re) for each company in the table. Which of the following statements is most likely TRUE: A, a 2 yer US treasury is yielding 2% and the expected market return is 1 196 B. a 2 yer US treasury is yielding 3% and the expected market return is 13% C. a 10 year US treosury is yielding 396 and the expected market return is 9% D, a 10 year US treasury is yielding 2% and the expected market return is 10% E, a 10 year US treasury is yielding 3% and the expected market return is 12% 3. This question has intentionally been left blank EVERYBODY GETS CREDIT A. B. C. D. E. 4. Company Alpha is also in the Widget industry. Currently, the company has $720M in total assets on the balance sheet and no debt. Thus, the yield to maturity (YTM) on existing debt does not exist. The CFO is considering adding up $360M of long term debt to the capital structure. Using the table of eomparable companies, the cost of debt for Company Alpha is approximately: A6% B, 896 С.996 D, 1196 E. 12% S. The company with the lowest weighted average cost of capital is: A. Company V B. Company W C. Company X D. Company Y E. Company Z 6. Based on the betas for all companies, the Widget industry i more likely non-cyclical than cyclica. Evaluate the underlined words in italies. True or False? A. True B. False
Answer #1

Answer1

Compared to larger company in widget industry,the smaller companies have lower effective tax rate

As shown in table company v and z have lower effective tax rate.

2.

Ist statement

as per capm= Rf+ Beta( Rm-Rf)=Re

A.

14.3= 2+1.25(11-2)=13.25 false

E.

15.6= 3+1.4(9)=15.6 true.

Therefore E is most likely true.

4.

As company is introducing 50% of assets as debt

it is approximately 6%

Answer 6

wideget industry is more likely to be non cyclical is true

because Beta 1= comapanies stock moves exactly with market

Beta greater than 1 means stock are volatile and non cyclic.